# SS Linear Fusion Model: Hyperspectral Imaging Efficient Spatial and Spectral Linear Model with Bidirectional Feature Learning

### v1

@@ IV-C Experimental Results and Analysis
This section presents empirical results from the SS Linear Model, evaluated on three widely-used hyperspectral datasets: Houston 2013, Indian Pines, and Pavia University. We report key metrics—Overall Accuracy (OA) and Kappa coefficient($\kappa$) )—and visually examine classification maps to compare the SS Linear Model’s classification accuracy with benchmark models.
@@ IV-C1 Houston 2013 Data Set
The Houston 2013 dataset is a challenging testbed due to its urban setting, diverse land-cover classes, and spectral variability. This dataset is ideal for evaluating the SS Linear Model’s ability to handle both spectral and spatial complexities. Table IV presents classification results, comparing the SS Linear Model to several state-of-the-art competitors, all evaluated with a standard patch size of 15, except for SS Linear Model’s Method 2, which uses a smaller patch size of 3.
@@
Performance Overview: SS Linear Model achieved impressive results, with Method 1 (patch size 15) closely trailing the best-performing model, HiT, by a small margin in terms of Overall Accuracy (OA). However, Method 2, which uses a patch size of 3, surpasses HiT by approximately 0.5% in OA, demonstrating that our model can achieve competitive and even superior performance with significantly less computational load. This efficiency showcases the model’s ability to achieve high accuracy even when operating at a smaller spatial context, which is often challenging in hyperspectral classification.
Efficiency and Resource Savings: Method 2’s success with a reduced patch size indicates that SS Linear Model is capable of capturing essential spatial-spectral features with smaller patches, thus reducing computational demands. By using a smaller patch size, Method 2 effectively lowers memory requirements and speeds up training and inference times, making it a resource-efficient choice for large-scale or real-time hyperspectral image analysis.
@@
Comparison with CNN and Transformer Models: Although transformers generally perform well in long-range spectral dependency modeling, they typically require large patches and high computational resources. The SS Linear Model, particularly in Method 2, balances spectral-spatial feature extraction with efficient processing, showing an advantage over CNNs in spectral detail capture and over transformers in computational efficiency.
The performance of Method 2 on the Houston 2013 dataset establishes the SS Linear Model as an efficient and accurate framework for hyperspectral classification, achieving state-of-the-art results with reduced patch size and computational resources. This makes it a suitable choice for applications where memory and speed are crucial constraints, without sacrificing classification precision.
Visualization and Qualitative Analysis and Comparison: Fig.2 provides a visual comparison of classification maps generated by SS Linear Model and other competing methods. The SS Linear Model’s classification map for Method 1 displays clean, well-defined boundaries between urban classes, with reduced misclassification in mixed-use areas such as Road and Railway. Compared to the results from larger patch sizes in other models, Method 1’s smaller patch size offers comparable or improved boundary precision, underscoring its effectiveness even with a reduced spatial context.
Figure 2: classification maps for the Houston 2013 dataset. Ground-Truth map, 9 comparative methods and Our SS Linear Model Method
@@ IV-C2 Indian Pines Data Set
The Indian Pines dataset presents unique challenges due to its mix of agricultural and forested land-cover classes with high spectral similarity. This environment tests the model’s ability to accurately separate classes with subtle spectral differences. Table V summarizes the classification results, where both SS Linear Model methods (Method 1 with patch size 15 and Method 2 with patch size 5) demonstrate strong performance.
@@
Performance Overview: SS Linear Model achieved an Overall Accuracy (OA) of 92.41% with Method 2 (patch size 5), slightly surpassing Method 1’s OA of 92.18% (patch size 15). Its performance is higher than 5% of the other competitors. This result not only positions Method 2 as the top-performing approach on this dataset but also underscores the model’s ability to maintain high classification accuracy with a reduced patch size. This is particularly advantageous for hyperspectral datasets like Indian Pines, where high spectral similarity across classes demands precise feature extraction.
@@
Class-wise Accuracy and Spectral Sensitivity: The SS Linear Model effectively handles the spectral complexity of agricultural classes, such as Corn-notill and Soybean-notill, which often have overlapping spectral signatures. Both methods demonstrate high Overall Accuracy (OA), with Method 1 and Method 2 showing consistent performance across most classes. For example, Method 1 achieves 100% accuracy in several challenging classes, such as Grass-pasture-mowed and Oats, which are traditionally difficult for many models due to their spectral resemblance to neighbouring vegetation classes.
@@
Efficiency with Smaller Patch Size: Method 2’s use of a patch size of 5, as opposed to the standard size of 15, highlights the SS Linear Model’s ability to balance computational efficiency with classification precision. By utilizing a smaller patch size, Method 2 reduces the memory footprint and speeds up training and inference times while still capturing essential spectral-spatial information. This efficiency is particularly beneficial in hyperspectral analysis, where data volumes are substantial and computational resources are often limited.
@@
Comparison with Transformer and CNN Models: Compared to traditional CNNs, which often underperform in capturing nuanced spectral differences across similar classes, SS Linear Model’s bidirectional spectral processing enhances its ability to handle subtle class distinctions. While transformer models are generally adept at long-range dependency modeling, they typically require larger patch sizes and higher computational resources. In contrast, Method 2’s smaller patch size and competitive performance highlight SS Linear Model as an efficient alternative for hyperspectral datasets like Indian Pines.
@@
Visualization and Qualitative Analysis and Comparison: Fig.3 illustrates the classification maps generated by the SS Linear Model and competing models. The map for Method 1 exhibits fewer misclassified pixels and clearer class boundaries, particularly in areas with mixed vegetation, such as Soybean-mintill and Grass-pasture. This visual clarity aligns with Method 1’s high accuracy and further demonstrates its ability to separate classes with subtle spectral differences, even with a smaller spatial context.
@@
Figure 3: Visualization and classification maps for the Indian Pines dataset. Ground-Truth map, 9 comparative methods and Our SS Linear Model Method
@@
The SS Linear Model’s performance on the Indian Pines dataset confirms its effectiveness in complex, spectrally similar environments, delivering both high accuracy and computational efficiency. Method 1’s ability to outperform larger patch-size models with a reduced spatial context makes it a practical choice for hyperspectral classification tasks where resources are constrained.
@@ IV-C3 University of Pavia Data Set
The University of Pavia dataset, an urban hyperspectral dataset with high spatial resolution, presents a unique challenge due to its detailed urban land-cover classes and high intra-class variability. This dataset is ideal for evaluating the model’s spatial-spectral processing capabilities. Table VI shows the classification results, where SS Linear Model, specifically Method 2 (patch size 5), demonstrates a good performance an close to the the big patch size 15 applied to these competitors.
@@
Performance Overview: On this dataset, Method 2 achieves an Overall Accuracy (OA) of 98.14%, surpassing most competing models, including the HiT transformer model, despite using a smaller patch size. Method 1, with a patch size of 15, also performs exceptionally well, reaching an OA of 96.35%. The high OA across both methods highlights SS Linear Model’s adaptability to urban landscapes with varied spatial and spectral characteristics.
@@
Resource Efficiency with Smaller Patch Size: Method 2’s use of a reduced patch size (3) not only results in competitive accuracy but also significantly lowers computational demands, reducing memory consumption and improving processing speed. This efficiency makes the SS Linear Model a practical solution for urban classification tasks where high-resolution data requires intensive computation. Method 2’s ability to maintain high accuracy while operating with a smaller spatial context demonstrates the model’s effective use of bidirectional spectral processing and spatial feature integration.
@@
Class-wise Accuracy and Robustness: Both Method 1 and Method 2 achieve high accuracy across challenging classes like Asphalt and Meadows, which are prone to spectral overlap with other urban features. Method 2, in particular, excels in distinguishing between spectrally similar classes, reaching over 99% accuracy in classes such as Metal Sheets and Bare Soil. This strong class-wise performance reflects SS Linear Model’s ability to capture fine-grained spatial-spectral details, a critical capability for accurate urban classification.
@@
Comparison with CNN and Transformer Models: Transformer-based models like HiT generally perform well in urban settings due to their ability to capture long-range dependencies. However, they require larger patches and substantial computational resources. SS Linear Model, particularly Method 2, offers a resource-efficient alternative that does not sacrifice accuracy, achieving comparable or superior results with a smaller patch size. Traditional CNNs, while effective for local spatial feature extraction, lack the spectral depth required for complex urban classification, where SS Linear Model’s bidirectional spectral processing offers a clear advantage.
@@
Visualization and Qualitative Analysis and Comparison: Figure 4 illustrates the classification maps generated by the SS Linear Model and competing models. Our method 1’s map shows well-defined boundaries and fewer misclassifications in areas like Gravel and Shadows, where traditional models often struggle. The clear separation of classes in Method 1’s map indicates its ability to handle intricate spatial details and maintain accuracy even with reduced patch sizes. This visual evidence aligns with Our method 1’s high quantitative performance, reinforcing its suitability for high-resolution urban data.
@@
Figure 4: Visualization and classification maps for the University Pavia dataset. Ground-Truth map, 9 comparative methods and Our SS Linear ModelMethod
@@
The results from the University of Pavia dataset confirm SS Linear Model’s effectiveness in urban classification, delivering both high accuracy and computational efficiency. Our Method’s performance highlights the model’s potential for applications requiring fast processing and resource efficiency without compromising classification quality, making it a strong candidate for real-time or large-scale urban hyperspectral analysis.
@@ IV-D Model Analysis
To better understand the inner workings and effectiveness of the SS Linear Model, we perform a comprehensive model analysis. This includes an ablation study to evaluate the impact of core components and a parameter sensitivity analysis focusing on patch size, which is critical for balancing accuracy and computational efficiency.
@@ IV-D1 Parameter Sensitivity Analysis
In hyperspectral image classification, patch size is a key parameter affecting both classification accuracy and computational efficiency. We conduct a parameter sensitivity analysis on the Houston 2013 dataset to evaluate the SS Linear Model’s performance across a range of patch sizes. Table VII summarizes the classification metrics across different patch sizes, with Figure 5 providing a visual comparison of Overall Accuracy (OA) trends.
Optimal Patch Size Determination: The results show that a patch size of 5 yields the highest accuracy for the SS Linear Model, establishing it as the optimal choice for balancing accuracy with computational demands on the Houston 2013 dataset. Smaller patch sizes (e.g., 3) slightly reduce OA but significantly decrease memory usage and training time, highlighting Method 2’s suitability for resource-constrained environments.
@@
Impact of Patch Size on Spatial-Spectral Resolution: Smaller patch sizes, such as 3 or 5, allow SS Linear Model to focus on fine-grained spectral details without overburdening computational resources. Larger patches provide more contextual information but are less efficient and may introduce redundant information in uniform areas. Method 2’s performance with patch size 5 suggests that SS Linear Model can effectively capture spatial-spectral patterns without needing extensive contextual input, validating its design for efficient hyperspectral classification.
@@
The ideal patch size varies by dataset. For instance, Houston and University of Pavia benefit from smaller patches (P3) due to their urban content and detailed spatial features. Meanwhile, Indian Pines, which contains more spectrally similar vegetation classes, achieves peak accuracy with intermediate patch sizes (P7-P11). This variability highlights the SS Linear Model’s adaptability and the importance of tuning patch size based on the spatial and spectral complexity of the dataset.
@@
Across datasets, the Kappa coefficient remains stable for smaller patches, indicating that SS Linear Model maintains class consistency even with reduced spatial context. This is particularly valuable for tasks requiring high precision in class-specific predictions. Smaller patches (P3) are recommended, particularly for real-time or mobile platforms where computational power is limited. Method 2, which uses P3 across all datasets, achieves near-peak accuracy and high Kappa values with reduced processing requirements. For detailed analysis or high-stakes applications (e.g., urban planning), larger patch sizes (P13 or P15) may offer slight accuracy advantages, especially in datasets like Pavia University with distinct urban classes.
@@ IV-D2 Ablation Study
The ablation study systematically examines the contribution of each major component in the SS Linear Model architecture: forward spectral processing, backward spectral processing, and spatial feature processing. By selectively removing these components and evaluating their impact on classification accuracy, we aim to clarify their roles in enhancing model performance. All ablation experiments were conducted on the Houston 2013 dataset with a consistent patch size of 3 across configurations to ensure comparability. The results are summarized in Table X.
Component Contribution Analysis: The complete SS Linear Model incorporates all three key components: forward spectral processing, backward spectral processing, and spatial feature processing. This full configuration achieves an Overall Accuracy (OA) of 97.60% and a Kappa coefficient of 97.40%, setting a baseline for comparison. As components are progressively removed, we observe notable declines in accuracy, underscoring the importance of each feature in the architecture.
Spatial Processing: When the spatial processing block is excluded (Method 2), the model’s OA drops to 94.73% and the Kappa score decreases to 94.31%. This reduction in accuracy highlights the importance of spatial information in hyperspectral classification, particularly in complex environments like urban areas, where spatial boundaries between classes are subtle. The spatial processing block allows the SS Linear Model to capture spatial dependencies across adjacent pixels, enabling it to better differentiate spatially proximate classes.
Forward and Backward Spectral Processing: Both forward and backward spectral processing pathways contribute to the model’s spectral feature extraction. Removing the forward pathway (Method 3) or the backward pathway (Method 4) results in a 93.78% and 93.37% reduction in OA, respectively. This decrease in performance suggests that unidirectional spectral processing limits the model’s ability to capture the full spectral context, highlighting the value of bidirectional processing. By integrating both directions, the SS Linear Model effectively encodes diverse spectral dependencies, which is essential for accurately classifying hyperspectral data.
Bidirectional Spectral Processing: The SS Linear Model’s bidirectional spectral processing uniquely positions it to capture comprehensive spectral dependencies. With forward and backward pathways, the model encodes a fuller spectral representation, enhancing its robustness against spectral variability in hyperspectral data. In this study, removing either direction leads to a classification accuracy drop of 2-3% on average, confirming that bidirectional processing provides a richer and more nuanced spectral representation than unidirectional approaches used in conventional CNN and transformer-based architectures.

### v2

@@ IV-C Experimental Results and Analysis
This section presents empirical results from the SS non-Linear Model, evaluated on three widely-used hyperspectral datasets: Houston 2013, Indian Pines, and Pavia University. We report key metrics—Overall Accuracy (OA) and Kappa coefficient($\kappa$) )—and visually examine classification maps to compare the SS non-Linear Model’s classification accuracy with benchmark models.
@@ IV-C1 Houston 2013 Data Set
The Houston 2013 dataset is a challenging testbed due to its urban setting, diverse land-cover classes, and spectral variability. This dataset is ideal for evaluating the SS non-Linear Model’s ability to handle both spectral and spatial complexities. Table IV presents classification results, comparing the SS non-Linear Model to several state-of-the-art competitors, all evaluated with a standard patch size of 15, except for SS non-Linear Model’s Method 2, which uses a smaller patch size of 3.
@@
Performance Overview: SS non-Linear Model achieved impressive results, with Method 1 (patch size 15) closely trailing the best-performing model, HiT, by a small margin in terms of Overall Accuracy (OA). However, Method 2, which uses a patch size of 3, surpasses HiT by approximately 0.5% in OA, demonstrating that our model can achieve competitive and even superior performance with significantly less computational load. This efficiency showcases the model’s ability to achieve high accuracy even when operating at a smaller spatial context, which is often challenging in hyperspectral classification.
Efficiency and Resource Savings: Method 2’s success with a reduced patch size indicates that SS non-Linear Model is capable of capturing essential spatial-spectral features with smaller patches, thus reducing computational demands. By using a smaller patch size, Method 2 effectively lowers memory requirements and speeds up training and inference times, making it a resource-efficient choice for large-scale or real-time hyperspectral image analysis.
@@
Comparison with CNN and Transformer Models: Although transformers generally perform well in long-range spectral dependency modeling, they typically require large patches and high computational resources. The SS non-Linear Model, particularly in Method 2, balances spectral-spatial feature extraction with efficient processing, showing an advantage over CNNs in spectral detail capture and over transformers in computational efficiency.
The performance of Method 2 on the Houston 2013 dataset establishes the SS non-Linear Model as an efficient and accurate framework for hyperspectral classification, achieving state-of-the-art results with reduced patch size and computational resources. This makes it a suitable choice for applications where memory and speed are crucial constraints, without sacrificing classification precision.
Visualization and Qualitative Analysis and Comparison: Fig.2 provides a visual comparison of classification maps generated by SS non-Linear Model and other competing methods. The SS non-Linear Model’s classification map for Method 1 displays clean, well-defined boundaries between urban classes, with reduced misclassification in mixed-use areas such as Road and Railway. Compared to the results from larger patch sizes in other models, Method 1’s smaller patch size offers comparable or improved boundary precision, underscoring its effectiveness even with a reduced spatial context.
Figure 2: classification maps for the Houston 2013 dataset. Ground-Truth map, 9 comparative methods and Our SS non-Linear Model Method
@@ IV-C2 Indian Pines Data Set
The Indian Pines dataset presents unique challenges due to its mix of agricultural and forested land-cover classes with high spectral similarity. This environment tests the model’s ability to accurately separate classes with subtle spectral differences. Table V summarizes the classification results, where both SS non-Linear Model methods (Method 1 with patch size 15 and Method 2 with patch size 5) demonstrate strong performance.
@@
Performance Overview: SS non-Linear Model achieved an Overall Accuracy (OA) of 92.41% with Method 2 (patch size 5), slightly surpassing Method 1’s OA of 92.18% (patch size 15). Its performance is higher than 5% of the other competitors. This result not only positions Method 2 as the top-performing approach on this dataset but also underscores the model’s ability to maintain high classification accuracy with a reduced patch size. This is particularly advantageous for hyperspectral datasets like Indian Pines, where high spectral similarity across classes demands precise feature extraction.
@@
Class-wise Accuracy and Spectral Sensitivity: The SS non-Linear Model effectively handles the spectral complexity of agricultural classes, such as Corn-notill and Soybean-notill, which often have overlapping spectral signatures. Both methods demonstrate high Overall Accuracy (OA), with Method 1 and Method 2 showing consistent performance across most classes. For example, Method 1 achieves 100% accuracy in several challenging classes, such as Grass-pasture-mowed and Oats, which are traditionally difficult for many models due to their spectral resemblance to neighbouring vegetation classes.
@@
Efficiency with Smaller Patch Size: Method 2’s use of a patch size of 5, as opposed to the standard size of 15, highlights the SS non-Linear Model’s ability to balance computational efficiency with classification precision. By utilizing a smaller patch size, Method 2 reduces the memory footprint and speeds up training and inference times while still capturing essential spectral-spatial information. This efficiency is particularly beneficial in hyperspectral analysis, where data volumes are substantial and computational resources are often limited.
@@
Comparison with Transformer and CNN Models: Compared to traditional CNNs, which often underperform in capturing nuanced spectral differences across similar classes, SS non-Linear Model’s bidirectional spectral processing enhances its ability to handle subtle class distinctions. While transformer models are generally adept at long-range dependency modeling, they typically require larger patch sizes and higher computational resources. In contrast, Method 2’s smaller patch size and competitive performance highlight SS non-Linear Model as an efficient alternative for hyperspectral datasets like Indian Pines.
@@
Visualization and Qualitative Analysis and Comparison: Fig.3 illustrates the classification maps generated by the SS non-Linear Model and competing models. The map for Method 1 exhibits fewer misclassified pixels and clearer class boundaries, particularly in areas with mixed vegetation, such as Soybean-mintill and Grass-pasture. This visual clarity aligns with Method 1’s high accuracy and further demonstrates its ability to separate classes with subtle spectral differences, even with a smaller spatial context.
@@
Figure 3: Visualization and classification maps for the Indian Pines dataset. Ground-Truth map, 9 comparative methods and Our SS non-Linear Model Method
@@
The SS non-Linear Model’s performance on the Indian Pines dataset confirms its effectiveness in complex, spectrally similar environments, delivering both high accuracy and computational efficiency. Method 1’s ability to outperform larger patch-size models with a reduced spatial context makes it a practical choice for hyperspectral classification tasks where resources are constrained.
@@ IV-C3 University of Pavia Data Set
The University of Pavia dataset, an urban hyperspectral dataset with high spatial resolution, presents a unique challenge due to its detailed urban land-cover classes and high intra-class variability. This dataset is ideal for evaluating the model’s spatial-spectral processing capabilities. Table VI shows the classification results, where SS non-Linear Model, specifically Method 2 (patch size 5), demonstrates a good performance an close to the the big patch size 15 applied to these competitors.
@@
Performance Overview: On this dataset, Method 2 achieves an Overall Accuracy (OA) of 98.14%, surpassing most competing models, including the HiT transformer model, despite using a smaller patch size. Method 1, with a patch size of 15, also performs exceptionally well, reaching an OA of 96.35%. The high OA across both methods highlights SS non-Linear Model’s adaptability to urban landscapes with varied spatial and spectral characteristics.
@@
Resource Efficiency with Smaller Patch Size: Method 2’s use of a reduced patch size (3) not only results in competitive accuracy but also significantly lowers computational demands, reducing memory consumption and improving processing speed. This efficiency makes the SS non-Linear Model a practical solution for urban classification tasks where high-resolution data requires intensive computation. Method 2’s ability to maintain high accuracy while operating with a smaller spatial context demonstrates the model’s effective use of bidirectional spectral processing and spatial feature integration.
@@
Class-wise Accuracy and Robustness: Both Method 1 and Method 2 achieve high accuracy across challenging classes like Asphalt and Meadows, which are prone to spectral overlap with other urban features. Method 2, in particular, excels in distinguishing between spectrally similar classes, reaching over 99% accuracy in classes such as Metal Sheets and Bare Soil. This strong class-wise performance reflects SS non-Linear Model’s ability to capture fine-grained spatial-spectral details, a critical capability for accurate urban classification.
@@
Comparison with CNN and Transformer Models: Transformer-based models like HiT generally perform well in urban settings due to their ability to capture long-range dependencies. However, they require larger patches and substantial computational resources. SS non-Linear Model, particularly Method 2, offers a resource-efficient alternative that does not sacrifice accuracy, achieving comparable or superior results with a smaller patch size. Traditional CNNs, while effective for local spatial feature extraction, lack the spectral depth required for complex urban classification, where SS non-Linear Model’s bidirectional spectral processing offers a clear advantage.
@@
Visualization and Qualitative Analysis and Comparison: Figure 4 illustrates the classification maps generated by the SS non-Linear Model and competing models. Our method 1’s map shows well-defined boundaries and fewer misclassifications in areas like Gravel and Shadows, where traditional models often struggle. The clear separation of classes in Method 1’s map indicates its ability to handle intricate spatial details and maintain accuracy even with reduced patch sizes. This visual evidence aligns with Our method 1’s high quantitative performance, reinforcing its suitability for high-resolution urban data.
@@
Figure 4: Visualization and classification maps for the University Pavia dataset. Ground-Truth map, 9 comparative methods and Our SS non-Linear ModelMethod
@@
The results from the University of Pavia dataset confirm SS non-Linear Model’s effectiveness in urban classification, delivering both high accuracy and computational efficiency. Our Method’s performance highlights the model’s potential for applications requiring fast processing and resource efficiency without compromising classification quality, making it a strong candidate for real-time or large-scale urban hyperspectral analysis.
@@ IV-D Model Analysis
To better understand the inner workings and effectiveness of the SS non-Linear Model, we perform a comprehensive model analysis. This includes an ablation study to evaluate the impact of core components and a parameter sensitivity analysis focusing on patch size, which is critical for balancing accuracy and computational efficiency.
@@ IV-D1 Parameter Sensitivity Analysis
In hyperspectral image classification, patch size is a key parameter affecting both classification accuracy and computational efficiency. We conduct a parameter sensitivity analysis on the Houston 2013 dataset to evaluate the SS non-Linear Model’s performance across a range of patch sizes. Table VII summarizes the classification metrics across different patch sizes, with Figure 5 providing a visual comparison of Overall Accuracy (OA) trends.
Optimal Patch Size Determination: The results show that a patch size of 5 yields the highest accuracy for the SS non-Linear Model, establishing it as the optimal choice for balancing accuracy with computational demands on the Houston 2013 dataset. Smaller patch sizes (e.g., 3) slightly reduce OA but significantly decrease memory usage and training time, highlighting Method 2’s suitability for resource-constrained environments.
@@
Impact of Patch Size on Spatial-Spectral Resolution: Smaller patch sizes, such as 3 or 5, allow SS non-Linear Model to focus on fine-grained spectral details without overburdening computational resources. Larger patches provide more contextual information but are less efficient and may introduce redundant information in uniform areas. Method 2’s performance with patch size 5 suggests that SS non-Linear Model can effectively capture spatial-spectral patterns without needing extensive contextual input, validating its design for efficient hyperspectral classification.
@@
The ideal patch size varies by dataset. For instance, Houston and University of Pavia benefit from smaller patches (P3) due to their urban content and detailed spatial features. Meanwhile, Indian Pines, which contains more spectrally similar vegetation classes, achieves peak accuracy with intermediate patch sizes (P7-P11). This variability highlights the SS non-Linear Model’s adaptability and the importance of tuning patch size based on the spatial and spectral complexity of the dataset.
@@
Across datasets, the Kappa coefficient remains stable for smaller patches, indicating that SS non-Linear Model maintains class consistency even with reduced spatial context. This is particularly valuable for tasks requiring high precision in class-specific predictions. Smaller patches (P3) are recommended, particularly for real-time or mobile platforms where computational power is limited. Method 2, which uses P3 across all datasets, achieves near-peak accuracy and high Kappa values with reduced processing requirements. For detailed analysis or high-stakes applications (e.g., urban planning), larger patch sizes (P13 or P15) may offer slight accuracy advantages, especially in datasets like Pavia University with distinct urban classes.
@@ IV-D2 Ablation Study
The ablation study systematically examines the contribution of each major component in the SS non-Linear Model architecture: forward spectral processing, backward spectral processing, and spatial feature processing. By selectively removing these components and evaluating their impact on classification accuracy, we aim to clarify their roles in enhancing model performance. All ablation experiments were conducted on the Houston 2013 dataset with a consistent patch size of 3 across configurations to ensure comparability. The results are summarized in Table X.
Component Contribution Analysis: The complete SS non-Linear Model incorporates all three key components: forward spectral processing, backward spectral processing, and spatial feature processing. This full configuration achieves an Overall Accuracy (OA) of 97.60% and a Kappa coefficient of 97.40%, setting a baseline for comparison. As components are progressively removed, we observe notable declines in accuracy, underscoring the importance of each feature in the architecture.
Spatial Processing: When the spatial processing block is excluded (Method 2), the model’s OA drops to 94.73% and the Kappa score decreases to 94.31%. This reduction in accuracy highlights the importance of spatial information in hyperspectral classification, particularly in complex environments like urban areas, where spatial boundaries between classes are subtle. The spatial processing block allows the SS non-Linear Model to capture spatial dependencies across adjacent pixels, enabling it to better differentiate spatially proximate classes.
Forward and Backward Spectral Processing: Both forward and backward spectral processing pathways contribute to the model’s spectral feature extraction. Removing the forward pathway (Method 3) or the backward pathway (Method 4) results in a 93.78% and 93.37% reduction in OA, respectively. This decrease in performance suggests that unidirectional spectral processing limits the model’s ability to capture the full spectral context, highlighting the value of bidirectional processing. By integrating both directions, the SS non-Linear Model effectively encodes diverse spectral dependencies, which is essential for accurately classifying hyperspectral data.
Bidirectional Spectral Processing: The SS non-Linear Model’s bidirectional spectral processing uniquely positions it to capture comprehensive spectral dependencies. With forward and backward pathways, the model encodes a fuller spectral representation, enhancing its robustness against spectral variability in hyperspectral data. In this study, removing either direction leads to a classification accuracy drop of 2-3% on average, confirming that bidirectional processing provides a richer and more nuanced spectral representation than unidirectional approaches used in conventional CNN and transformer-based architectures.